### v2

# IR Bounds on Theories with Spontaneously-Broken Lorentz Symmetry

Abstract
In nature, some UV features of the dynamics are reflected in IR quantities. In fully relativistic theories, this connection can be probed through the analyticity properties of scattering amplitudes, allowing to understand which IR theories respect the UV assumptions of quantum field theory. The ensuing analyticity bounds can be usually rephrased as the absence of faster-than-light propagation for low-energy excitations. While it is interesting to understand these relations and their IR characterization for theories that have less idealized properties, it is also more difficult to derive analyticity bounds in these cases. For theories that spontaneously break Lorentz symmetry, recent progress was made by considering correlators of conserved currents and their analyticity properties. In this work, we focus on such theories and close the gap from the IR side, deriving bounds from the speed of propagation that are equivalent to the known analyticity bounds. Our bounds require that gapped excitations have a slower speed than gapless ones, at least for momenta that are low with respect to the mass gap. These results suggest a way to study the UV/IR connection in more complex theories.
1 Introduction
Separation of scales in nature allows to describe low-energy (IR) dynamics without knowledge of high-energy (UV) degrees of freedom. At the same time, low-energy interactions encode some of the properties that characterize the high-energy dynamics.
This connection between UV and IR becomes remarkably transparent for theories that in the UV are compatible with unitarity, Lorentz invariance, locality and microcausality – the property that operators commute when space-like separated Eden1971 . These assumptions about the UV theory entail polynomial boundedness and canonical analyticity properties of S-matrix elements as functions of the Mandelstam invariants Bros1964 ; Martin:1969 ; Streater:1989 , leading to bounds on the coefficients of low-energy interactions, see e.g. Pham:1985 ; Adams:2006sv ; deRham:2017avq ; Caron-Huot:2020cmc ; Bellazzini:2020cot ; Tolley2021 ; Sinha:2020win ; Bellazzini2022 ; Arkani-Hamed2021a ; Caron-Huot2022 .
In light of this UV/IR connection, it is interesting to explore whether these boundaries in the space of low-energy Wilson coefficients can be characterized in a synthetic way in terms of IR dynamical features. Such an IR characterization has been explored e.g. in Adams:2006sv ; Camanho:2014apa ; deRham:2021bll ; Serra2022 ; CarrilloGonzalez2022 ; CarrilloGonzalez2023 ; Serra:2023nrn , finding that certain effective field theories (EFTs) that do not respect the analyticity bounds display faster-than-light propagation for excitations around non-trivial backgrounds. These results exposed that the phase velocity captures, at least in some cases, the UV properties that lead to analyticity bounds. In particular, in various examples of Lorentz invariant theories, one can argue (and sometimes prove) that analyticity bounds forbid classical time advances at least on a wide class of backgrounds; see e.g. Arkani-Hamed2021a ; Caron-Huot2022 ; Bellazzini:2022wzv .
This may be phrased in terms of the relation between microcausality and phase velocity. Indeed, the limiting value of the phase velocity at high frequency, a.k.a. the front velocity, should be insensitive to any fixed wavelength background and should therefore fall into the light-cone as dictated by microcausality Brillouin:1960tos ; Diener:1996mj . Therefore, for cases in which the phase velocity does not decrease in the UV, its value at low frequency should always be lower than the speed determined by microcausality [^1] See e.g., Diener:1996mj ; Greene:2022uyf for interesting exceptions in this regard..
Given these results, it is valuable to understand how the same connection between UV and IR becomes manifest in more general contexts, where Lorentz invariance, locality, or other assumptions may be to some extent relaxed, in either or both the UV and the IR. For instance, EFTs describing physical systems like condensed matter or cosmology may carry information about the underlying UV theory, even if the low-energy interactions nonlinearly realize Lorentz. As a matter of fact, analyticity bounds are rooted in the study of propagation of light in a material, see e.g. Keldysh1989TheDF ; Creminelli:2024lhd and have recently been fruitfully extended to more generic theories that spontaneously break Lorentz invariance Creminelli:2022onn . Similarly, less conservative assumptions about the UV may still entail rich analyticity structures and map out boundaries in the space of Wilson coefficients that might be measured in the future, see e.g. Haring:2022cyf ; Buoninfante:2023dyd ; Buoninfante:2024ibt for recent efforts in this direction. In this regard, it would be very interesting to understand whether analyticity bounds and the corresponding UV assumptions have a universal translation in terms of IR dynamical features.
In this work, as a first step towards understanding the IR characterization of analyticity bounds beyond the most conservative assumptions, we focus on the case of an EFT that spontaneously breaks Lorentz symmetry. In this context the analyticity of correlators of conserved currents can be exploited to obtain bounds on the coefficients of low-energy interactions Creminelli:2022onn . We parallel these results by studying how the speed of propagation of low-energy excitations in the EFT is affected by interactions, deriving bounds on the Wilson coefficients in terms of the excitations’ velocity. In this way, we provide the IR analog of the analyticity bounds derived in Creminelli:2022onn , confirming that the velocity of low-energy excitations can meaningfully capture the information contained in the analyticity bounds. As we discuss below, we find that a more general notion of velocity, going beyond the phase velocity, is needed to characterize the IR system in a synthetic way.
In the specific, we consider the EFT of a conformal superfluid coupled to a gauge field in $D=3$ space-time dimensions. These systems are characterized by a non-linear realization of time translations due to a finite charge density phase of a scalar field charged under $U(1)$. This phase breaks the conformal symmetry to spatial rotations and translations plus the non-linearly realized time-translation. The IR dynamics of these systems has been studied extensively in the context of large charge expansion, see e.g. Hellerman:2015nra ; Monin:2016jmo ; Cuomo:2020rgt ; Cuomo:2021qws .
In these systems, the existence and analyticity properties of the S-matrix become nontrivial Hui:2023pxc ; Creminelli:2023kze , requiring one to consider alternative approaches in order to derive analyticity bounds. In Creminelli:2022onn this was done by studying correlators of the conserved $U(1)$ currents, due to their convenient UV behavior and analytic properties. In practice, these correlators were computed by adding a background, non-dynamical $U(1)$ gauge field coupled to the superfluid. While the $U(1)$ symmetry can be easily gauged, it is not simple to study the dynamics of the gauge field and to prospect which UV completions would give rise to the gauged EFT in the broken phase.
In our work, we follow an approach similar in spirit to what was done in Creminelli:2022onn and introduce a $U(1)$ gauge field coupled to the conformal superfluid, with vanishing background field-strength. Differently from Creminelli:2022onn , we consider the gauge field in this IR system as dynamical. While it might be subtle to interpret this IR system as a physical system that is UV completed by a conformal field theory, we find it useful to consider this setup simply as an auxiliary system to extract key information about the conformal superfluid, similarly to how the background gauge field is used in Creminelli:2022onn . In this IR system, we imagine that a background charge density makes the configuration with vanishing background field strength stable. Our point of view follows the natural expectation that a background field may be regarded as an approximation of a dynamical field whose excitations are being neglected. As a matter of fact, the quadratic part of the effective action for the $U(1)$-gauged conformal superfluid allows to consistently describe the excitations of both the scalar and vector fields, even when one writes this action imagining a non-dynamical, background vector field. In this action, a kinetic mixing converts scalar modes into vector modes and viceversa. As we discuss, this kinetic mixing allows to recover the information contained in correlators of conserved currents by examining the propagation of the low-energy modes.
A universal feature of the EFTs for the broken phase is that gapless excitations at zero momentum have soundspeed $c_{s}^{2}=1/2$ in $D=3$. This means that in any UV complete model that has Lorentz symmetry in the UV, the phase velocity of a gapless mode will change from $1/2$ to $1$ as one explores excitations with higher and higher momenta (eventually outside of the regime of validity of the EFT for the broken phase). If the low-energy theory has a mass gap, the phase velocity cannot be used to characterize the system at momenta below the mass gap, as it is divergent at zero momentum. A better notion of velocity is the one associated with the trajectories of rays Boillat:1969slt ; Boillat:1970gw ; Landau ; Sawicki:2024ryt , which coincides with the phase velocity for gapless excitations and has a meaningful non-relativistic limit (vanishing at zero momentum, for gapped excitations). With this, one can formulate a meaningful bound for the EFT by asking that the ray velocity of gapped excitations does not grow too much for momenta below the mass gap. We apply this reasoning to the gauged conformal superfluid, since the kinetic mixing between scalar and vector excitations leads to a mass gap for both scalar and vector modes. In practice we ask the ray velocity to be at most $1/2$ for momenta lower than the mass gap. In this way, the excursion between IR and UV values of ray velocity is entirely captured by the (effectively) gapless, high momentum regime of the theory, as one might expect in analogy with the gapless cases. This IR requirement, as we show, leads to the same bounds found in Creminelli:2022onn studying correlators of conserved currents.
This result makes it clear that the IR meaning of the bounds is not the absence of paradoxical causality violations, as previously thought Adams:2006sv . Indeed, propagation outside the speed $1/2$ soundcone is not in conflict with causality, while it is forbidden by the analyticity bounds. This is consistent with the finding that no causal paradoxes associated with superluminality in the IR can be predicted in the EFT, as discussed in Babichev:2007dw ; Burrage:2011cr ; Kaplan:2024qtf . The work is organized as follows. In Section 2 we describe the model of gauged conformal superfluid and its propagating modes at low frequencies. In Section 3 we present a precise derivation of ray velocity for gapless and gapped modes using the acoustic metric description. In Section 4, we obtain the IR bounds on the EFT of conformal superfluids using the ray velocity, and compare our result with the analyticity bounds of Creminelli:2022onn . We close in Section 5 with a brief discussion. Throughout this paper we use the (,+,,+)(-,+$\dots$,+)( - , + , … , + ) signature convention.
2 Conformal superfluid coupled to a gauge field
We start by considering an effective field theory for a conformal superfluid in $D=3$ coupled to a vector field, as discussed in Ref. Creminelli:2022onn ,
$\mathcal{L}$ $=$ $\frac{c_{1}}{6}|\nabla\chi|^{3}-2c_{2}\frac{(\partial|\nabla\chi|)^{2}}{|\nabla\chi|}+c_{3}\left(2\frac{(\nabla^{\mu}\chi\partial_{\mu}|\nabla\chi|)^{2}}{|\nabla\chi|^{3}}+\partial_{\mu}\left(\frac{\nabla^{\mu}\chi\nabla^{\nu}\chi}{|\nabla\chi|^{2}}\right)\partial_{\nu}|\nabla\chi|\right)$ (1)
$-\frac{b}{4}\frac{F_{\mu\nu}F^{\mu\nu}}{|\nabla\chi|}+\frac{d}{2}\frac{F^{\mu}{}_{i}F^{\nu i}}{|\nabla\chi|^{3}}\nabla_{\mu}\chi\nabla_{\nu}\chi\,,$
where
$\nabla_{\mu}\chi$$\equiv$ $\partial_{\mu}\chi-A_{\mu}\,,$ (2)
$|v|$$\equiv$ $\sqrt{-v_{\mu}v^{\mu}}\,.$ (3)
As we outline in Sec. 4.3, such an EFT arises e.g. when integrating out the radial mode of a $U(1)$ charged scalar in a conformal field theory. If the vector field was a dynamical degree of freedom in the UV, then one would expect a purely kinetic term for the gauge field to appear in Eq. (1), for instance derived using the symmetry arguments in Cuomo:2021qws . To keep clear the analogy with the work in Creminelli:2022onn , we will neglect this term and use the above EFT only as an auxiliary IR system to extract information about the conformal superfluid. As we show below, the interactions between scalar and vector already endow the vector with a healthy kinetic term in the broken phase. This theory only makes sense on top of a time-dependent background for the scalar field
$\chi(x)=\mu t+\pi(x)\,,$ (4)
which is responsible for the spontaneous breaking of Lorentz invariance below the scale $\mu$, which we assume $\mu>0$ without loss of generality, and which acts as the cutoff of the EFT for the broken phase. This theory enjoys a gauge symmetry
$\pi\to\pi+\Lambda\qquad;\qquad A_{\mu}\to A_{\mu}+\partial_{\mu}\Lambda\,.$ (5)
In order to study the dynamics of the low-energy degrees of freedom in the broken phase we turn to the quadratic Lagrangian for both $\pi$ and $A_{\mu}$,
$\mathcal{L}_{2}$ $=$ $\frac{c_{1}\mu^{3}}{6}+\frac{\mu c_{1}}{2}\left[\left(\dot{\pi}+A^{0}\right)^{2}-\frac{1}{2}\left(\partial_{i}\pi-A_{i}\right)^{2}+\mu\left(\dot{\pi}+A^{0}\right)\right]+\frac{2c_{2}}{\mu}\left[-\pi\square\ddot{\pi}+2A^{0}\square\dot{\pi}+A^{0}\square A^{0}\right]$ (6)
$+\frac{2c_{3}}{\mu}\left[-\pi\square\ddot{\pi}+2A^{0}\square_{c_{s}}\dot{\pi}-A^{i}\partial_{i}\ddot{\pi}+(\dot{A}^{0})^{2}+\dot{A}^{0}\partial_{i}A^{i}\right]$
$+\frac{(b+d)}{2\mu}\left[(\partial_{i}A^{0})^{2}+(\partial_{0}A_{i})^{2}+2\dot{A}^{0}(\partial_{i}A_{i})\right]-\frac{b}{4\mu}(\partial_{i}A_{j}-\partial_{j}A_{i})^{2}\,,$
with $i=1,2$ running over the spatial directions. This Lagrangian provides both kinetic terms for $\pi$ and for $A_{\mu}$, which are free of ghost, gradient and tachyonic instabilities for $c_{1}>0$, $b>0$ and $b+d>0$. Therefore, this quadratic Lagrangian and the related Hamiltonian, consistently describe the vector and scalar fields as healthy, dynamical variables. Importantly, there is also kinetic mixing between $\pi$ and $A_{\mu}$ which implies that if one excites either of the two fields, inevitably the other will get excited as well. Due to this, we should study the eigenmodes of propagation.
As discussed in the Introduction, we do not insist that the gauge field should be a physical degree of freedom of the UV conformal theory, and content ourselves with regarding it as part of an IR auxiliary system we use to study the conformal superfluid. Auxiliary or not, we imagine that the IR system describing the broken phase with vanishing electric potential is stable thanks to a suitable homogeneous background charge. This background charge makes possible for the scalar field to get a non-zero charge density while satisfying the Gauss law with zero background electric field. This phase will be stable if one chooses appropriate self interactions for the scalar.
2.1 Scalar gapless mode
It is instructive to first consider the decoupling limit when the vector field is in fact nondynamical (formally $b+d\to\infty$). In this case, the dispersion relation for the canonical scalar $\pi_{c}=\sqrt{\mu c_{1}}\pi$ is obtained as
$G^{-1}_{\pi\pi}(\omega,{\bf p})$ $=$ $\left(\omega^{2}-\frac{{\bf p}^{2}}{2}\right)+\frac{4(c_{2}+c_{3})}{\mu^{2}c_{1}}\left(\omega^{2}-{\bf p}^{2}\right)\omega^{2}=0\,.$ (7)
When solved perturbatively in $c_{2}$ and $c_{3}$, this gives,
$\omega_{\pi}^{2}({\bf p})=\frac{{\bf p}^{2}}{2}+\frac{4(c_{2}+c_{3})}{\mu^{2}c_{1}}\left(\frac{{\bf p}^{2}}{2}\right)^{2}\,.$ (8)
As expected, due to the conformal symmetry in $D=3$ and the breaking of Lorentz invariance below $\mu$, the speed of propagation at zero momentum is $c_{s}^{2}=1/2$. In general we find a momentum dependent phase velocity [^2] In this case the usual definition holds since the background is spatially isotropic and there is no gap. A more general definition is provided in Sec. 3. $\omega_{\pi}^{2}({\bf p})/{\bf p}^{2}$, which either increases or decreases with respect to $1/2$ depending on the sign of $c_{2}+c_{3}$. At this point, we only notice that the phase velocity will increase monotonically for increasing (small) momenta only if $c_{2}+c_{3}>0$. However, there is no further bound we can impose at this stage on the size of these coefficients, other than the requirement of validity of the perturbative treatment within the EFT, which is rather uninformative.
2.2 Dynamical gauge field
As discussed before, in general for finite $b$ and $d$ the quadratic Lagrangian (6) describes not only a propagating scalar, but also a propagating gauge field. In fact, these are kinetically mixed and therefore the eigenmodes of propagation are a combination of both. Interestingly, this endows the would-be scalar mode with a gap.
We proceed to find the propagating degrees of freedom. The second order Lagrangian can be written, in Fourier space, as
$\mathcal{L}_{2}=\frac{1}{2}\Phi_{I}^{\dagger}(-\omega,-{\bf p})\,\mathcal{K}_{IJ}(\omega,{\bf p})\Phi_{J}(\omega,{\bf p})\,,$ (9)
where $\Phi_{I}=\left(\bar{\pi},\bar{A}^{\nu}\right)$, and $\bar{\pi}=\sqrt{\mu}\,\pi$ and $\bar{A}^{\nu}=\frac{1}{\sqrt{\mu}}A^{\nu}$ are the dimension-$1/2$ fields. The kinetic $4\times 4$ matrix $\mathcal{K}_{IJ}(\omega,{\bf p})$ is given by
$\mathcal{K}_{IJ}(\omega,{\bf p})=\begin{pmatrix}c_{1}G^{-1}_{\pi\pi}&i\hat{J}_{0}&i\hat{J}_{j}\\ -i\hat{J}_{0}&u&v\,p_{j}\\ -i\hat{J}_{i}&v\,p_{i}&V\delta_{ij}+Wp_{i}p_{j}\\ \end{pmatrix}\,,$ (10)
with
$\hat{J}_{0}$ $=$ $\left[\mu^{2}c_{1}+4c_{2}\left(\omega^{2}-{\bf p}^{2}\right)+4c_{3}\left(\omega^{2}-\frac{{\bf p}^{2}}{2}\right)\right]\frac{\omega}{\mu}\,,$ (11)
$\hat{J}_{i}$ $=$ $-\left(\frac{\mu^{2}c_{1}}{2}+2c_{3}\,\omega^{2}\right)\frac{p_{i}}{\mu}\,,$ (12)
$u$ $=$ $\mu^{2}c_{1}+4c_{2}\left(\omega^{2}-{\bf p}^{2}\right)+4c_{3}\,\omega^{2}+(b+d){\bf p}^{2}\,,$ (13)
$v$ $=$ $-\left(2c_{3}+(b+d)\right)\omega\,,$ (14)
$V$ $=$ $-\frac{\mu^{2}c_{1}}{2}+(b+d)\omega^{2}-b\,{\bf p}^{2}\,,$ (15)
$W$ $=$ $b\,.$ (16)
The dispersion relations $\omega^{2}({\bf p})$ for the propagating degrees of freedom are obtained from the condition $\det(\mathcal{K})=0$. To make sense of this procedure, we fix a “Coulomb”-type gauge $\partial_{i}A^{i}=0$. Other choices of gauge give the same result. We find the physical degrees of freedom to be determined by the following equations
00 0 $=$ $V\,,$ (17)
00 0 $=$ $u\left(V+W{\bf p}^{2}\right)-v^{2}{\bf p}^{2}\,.$ (18)
The first equation describes the transverse mode, with dispersion relation
$\omega_{\perp}^{2}({\bf p})$ $=$ $m^{2}+\frac{b}{(b+d)}{\bf p}^{2}\,,$ (19)
where
$m^{2}=\frac{\mu^{2}c_{1}}{2(b+d)}\,.$ (20)
Since there are no contributions from the higher-dimensional operators $c_{2}$ and $c_{3}$ in $\omega_{\perp}^{2}({\bf p})$, we will not examine this further. The second one is a mix of the scalar and $A^{0}$ modes (after removing the longitudinal mode by the gauge fixing condition). In this second dispersion relation the size of $c_{2}$ and $c_{3}$ controls contributions with higher powers in $\omega$ and/or ${\bf p}$. Due to this, Eq. (18) contains in fact both low-energy and high-energy modes. Focusing on the low-energy mode, we can solve perturbatively at linear order in $c_{2,3}$ within the regime of validity of the EFT. This leads to:
$\omega_{-}^{2}({\bf p})$ $=$ $m^{2}+\frac{{\bf p}^{2}}{2}-\frac{4(c_{2}+c_{3})}{\mu^{2}c_{1}}\left(-\frac{{\bf p}^{2}}{2}+m^{2}\right)\left(\frac{{\bf p}^{2}}{2}+m^{2}\right)$ (21)
$+\frac{\mu}{2(b+d)}\left[\frac{4c_{2}}{\mu}\left(-\frac{{\bf p}^{2}}{2}+m^{2}\right)+\frac{4c_{3}}{\mu}\left(\frac{{\bf p}^{2}}{2}+m^{2}\right)\right]\,.$
Notice that when $c_{1}\ll(b+d)$, the vector effectively decouples from the scalar and also $m^{2}\to 0$, thus recovering the gapless scalar mode from Eq. (8). Keeping instead $b\,,\,d$ finite, the above equation has the form:
$\omega_{-}^{2}({\bf p})$ $=$ $m^{2}+\alpha^{2}({\bf p})\,{\bf p}^{2}\,,$ (22)
where we have defined the momentum-dependent parameter $\alpha^{2}({\bf p})$ as
$\alpha^{2}({\bf p})=\frac{1}{2}+\tilde{c}_{2}\left(-1+\frac{{\bf p}^{2}}{2m^{2}}\right)+\tilde{c}_{3}\left(1+\frac{{\bf p}^{2}}{2m^{2}}\right)\,,$ (23)
with $\tilde{c}_{2,3}\equiv c_{2,3}/(b+d)$.
One would be tempted to use the coefficient of the ${\bf p}^{2}$ term in Eq. (22), namely $\alpha^{2}({\bf p})$, as a notion of velocity. However, as we will discuss in the next section, this requires some care due to the presence of a gap. After clarifying this, we will use the dispersion relation (22) to derive bounds in Sec. 4.
3 Ray velocity from the acoustic metric
In order to study the propagation of the low-energy modes in this theory, it is useful to turn to the effective description given by the acoustic metric. The purpose of this Section is to provide a general definition of velocity that can be used for gapped modes and arbitrary backgrounds, and which reduces to the usual phase velocity for gapless modes. This is the so-called ray velocity Boillat:1969slt ; Boillat:1970gw ; Landau , which is given in Eq. (32), and which will prove important for the formulation of our bounds. The reader interested only on the bounds may skip this section and instead go directly to Sec. 4. For a more detailed account of this formalism see Ref. Sawicki:2024ryt .
We begin by considering a theory with gapless excitations described by a quadratic Lagrangian of the form
$\mathcal{L}_{2}=-\frac{1}{2}\sqrt{-S}Z^{\mu\nu}\partial_{\mu}\delta\chi\partial_{\nu}\delta\chi\,,$ (24)
where $Z^{\mu\nu}$ is the acoustic metric and $S_{\mu\nu}$ its inverse, and $S=\det(S_{\mu\nu})$. In this case the associated linearized equation of motion (EoM) can be written as
$\frac{1}{\sqrt{-S}}\partial_{\mu}\left(\sqrt{-S}Z^{\mu\nu}\partial_{\nu}\delta\chi\right)=\bar{\square}\delta\chi=0\,,$ (25)
with ¯$\bar{\square}=Z^{\mu\nu}\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}$ and ¯$\bar{\nabla}_{\mu}$ is the covariant derivative compatible with $Z$ and $S$. This is simply the wave-equation on the nontrivial background given by the acoustic metric.
It is possible to study this equation in the so called Eikonal expansion, with an ansatz $\delta\chi=A(x)\,e^{i\mathcal{S}(x)/\epsilon}$, and $\epsilon\to 0$. This amounts to consider a rapidly-varying phase $\mathcal{S}(x)$ and a slowly-varying amplitude $A(x)$. One can show that at leading order in $1/\epsilon$, that is the geometric optics approximation, Eq. (25) corresponds to
$Z^{\mu\nu}p_{\mu}p_{\nu}=0\,,$ (26)
with $p_{\mu}=\partial_{\mu}\mathcal{S}$, which effectively picks up the terms in the linearized EoM with the most derivatives acting on the fluctuations. To properly regard $Z^{\mu\nu}$ as an effective metric, we should be able from the above to derive a corresponding geodesic equation. Acting with the covariant derivative ¯$\bar{\nabla}_{\alpha}$ on the above we get
$Z^{\mu\nu}p_{\nu}\bar{\nabla}_{\mu}p_{\alpha}=0\,,$ (27)
where we used the fact that $p_{\mu}$ is the gradient of a scalar quantity, the phase $\mathcal{S}$, and therefore ¯$\bar{\nabla}_{\alpha}p_{\mu}=\bar{\nabla}_{\mu}p_{\alpha}$. Finally, contracting with $Z^{\nu\alpha}$ and defining the “ray-vector” as
$N^{\mu}\equiv Z^{\mu\nu}p_{\nu}\,,$ (28)
we obtain the geodesic equation
$N^{\mu}\bar{\nabla}_{\mu}N^{\nu}=0\,.$ (29)
In light of this, Eq. (27) can be rephrased as the parallel transport of energy-momentum along the $N^{\mu}$ geodesic, $N^{\mu}\bar{\nabla}_{\mu}p_{\nu}=0$. Physically, the ray vector $N^{\mu}$ tracks the flow of energy-momentum $p_{\mu}$. It is also useful to use the definition of $N^{\mu}$ to rewrite the EoM as
$S_{\mu\nu}N^{\mu}N^{\nu}=0\,.$ (30)
In conclusion, the gapless EoM implies that the contravariant vector $N^{\mu}$ follows null geodesics of $S_{\mu\nu}$. This defines the trajectories, a.k.a “rays”, describing propagation of the wavefronts in the geometric optics approximation,
$N^{\mu}p_{\mu}=N^{\mu}\partial_{\mu}\mathcal{S}=0\quad,\quad\text{for gapless modes.}$. (31)
Since $N^{\mu}$ is the tangent vector of the rays, it is natural to define the speed of propagation along the geodesic – the ray velocity – in a given coordinate frame as
${\bf v}=\frac{{\bf N}}{N^{0}}\,.$ (32)
This is a general definition which is valid for arbitrary backgrounds, and can even accommodate the presence of a gap, as we will explain below in Sec. 3.2. In the case of a gapless mode, $N^{\mu}$ parametrizes the trajectories of the constant phase surfaces as expressed by Eq. (31), meaning that Eq. (32) coincides with the phase velocity of a gapless mode Landau ; Sawicki:2024ryt . Note that $N^{\mu}$ should be normalized to have dimensions of a length.
3.1 An example: P(X) theory
To see that this treatment reproduces the usual known results, let us turn to the simple example of a $P(X)$ theory with Lagrangian
$\mathcal{L}=-\frac{1}{2}(\partial\phi)^{2}+\frac{1}{4}c_{2}(\partial\phi)^{4}\,.$ (33)
The corresponding acoustic metric can be obtained by expanding at quadratic order in perturbations around a fixed background, $\phi=\bar{\phi}+\varphi$, and it reads
$Z^{\mu\nu}$ $=$ $\left(1-c_{2}(\partial{\phi})^{2}\right)\eta^{\mu\nu}-2c_{2}\,\partial^{\mu}{\phi}\,\partial^{\nu}{\phi}\,,$ (34)
which is generally spatially anisotropic except in the frame where ${\vec{\nabla}}\phi=0$, which only exists if $(\partial\phi)^{2}<0$. Here for ease of notation we dropped the bar from the background quantities. The corresponding dispersion relation is
$\omega^{2}({\bf p})={\bf p}^{2}-\frac{2c_{2}(\partial\phi\cdot p)^{2}}{(1-c_{2}(\partial\phi)^{2})}\,,$ (35)
which is gapless. We proceed to compute the inverse of $Z^{\mu\nu}$, which is easily found to be
$S_{\mu\nu}$ $=$ $\frac{1}{\left(1-c_{2}(\partial{\phi})^{2}\right)}\left[\eta_{\mu\nu}+\frac{2c_{2}\,\partial_{\mu}{\phi}\,\partial_{\nu}{\phi}}{\left(1-3c_{2}(\partial{\phi})^{2}\right)}\right]\,.$ (36)
which allows to write the EoM in the geometric optics approximation in terms of $N^{\mu}$, obtaining
$\eta_{\mu\nu}N^{\mu}N^{\nu}+\frac{2c_{2}\,(\partial{\phi}\cdot N)^{2}}{\left(1-3c_{2}(\partial{\phi})^{2}\right)}$ $=$ $0\,.$ (37)
From here we can compute the ray velocity as defined in Eq. (32),
${\bf v}^{2}=1-\frac{2c_{2}(\partial\phi\cdot N)^{2}}{(1-3c_{2}(\partial\phi)^{2})(N^{0})^{2}}\,,$ (38)
which coincides with the usual definition of phase velocity $\omega^{2}({\bf p})/{\bf p}^{2}$ for a background with spatially isotropic acoustic metric [^3] See the discussion in Sawicki:2024ryt . If we impose $N^{\mu}p_{\mu}=0$, then it follows that $\frac{{\bf N}\cdot{\bf p}}{N^{0}\omega}=1$, meaning that $\omega^{2}/{\bf p}^{2}={\bf N}^{2}/(N^{0})^{2}$, provided ${\bf p}$ is parallel to ${\bf N}$..
Suppose that in this theory we want to avoid superluminality. Requiring the ray velocity (38) to be less or equal to the speed of light $c=1$ immediately implies [^4] Here we are assuming that the denominator above is always positive, which is guaranteed by the absence-of-ghosts condition.
$c_{2}\geq 0\,.$ (39)
This is the well known condition to prevent superluminality in this model Adams:2006sv . Geometrically, this means the ray-vector must be timelike (or null) according to the standard spacetime metric $\eta_{\mu\nu}$, i.e. lie inside (or on) the lightcone. Using Eq. (37) above
$\eta_{\mu\nu}N^{\mu}N^{\nu}=-\frac{2c_{2}\,(\partial{\phi}\cdot N)^{2}}{\left(1-3c_{2}(\partial{\phi})^{2}\right)}\leq 0\,,$ (40)
which again implies (39).
As a closing remark, in this case due to the absence of a mass term, we could have also obtained this result by instead demanding that the momentum vector lies outside (or on) the lightcone, that is, it is $\eta$-spacelike (or null). From the EoM in terms of $p_{\mu}$ we get
$\eta^{\mu\nu}p_{\mu}p_{\nu}=\frac{2c_{2}\,(\partial{\phi}\cdot p)^{2}}{\left(1-c_{2}(\partial{\phi})^{2}\right)}\geq 0\,,$ (41)
again leading to (39). However, in general this does not imply that the ray velocity agrees with $\omega^{2}({\bf p})/{\bf p}^{2}$. This is especially important when comparing it against a $c_{s}^{2}\neq 1$ or when handling a theory with a mass gap.
3.2 Beyond the zero-gap approximation
For our particular purpose of studying the gapped mode (22) in the low-momentum regime, we must consider the opposite limit where momenta are not necessarily large enough to allow us to neglect the terms in the EoM with less derivatives. To keep the discussion simple we consider just a constant term [^5] This is justified in our case since we have a constant acoustic metric ($\partial_{\rho}Z^{\mu\nu}=0$, and therefore we can simply consider a plane-wave ansatz $\delta\chi=A\,e^{ip_{\mu}x^{\mu}}$, with $A$ and $p_{\mu}$ constants, instead of the more general Eikonal ansatz.,
$Z^{\mu\nu}p_{\mu}p_{\nu}+C^{2}=0\,,$ (42)
which implies a gap $m^{2}=C^{2}/|Z^{00}|$.
Following the same steps as before one can prove that $N^{\mu}=Z^{\mu\nu}p_{\nu}$ still follows the geodesic equation Eq. (29), but now the geodesics are $S$-timelike,
$S_{\mu\nu}N^{\mu}N^{\nu}=-C^{2}\,.$ (43)
For our purposes this is the only change with respect to the previous case where the gap was negligible. Therefore the geometrical interpretation remains the same and the phase velocity can be defined as before by Eq. (32). Importantly, as a consequence of the presence of a gap we have $N^{\mu}p_{\mu}\neq 0$, meaning that in this case the speed defined through $N^{\mu}$ is no longer the phase velocity.
As a short and rather trivial example, we can consider a Lorentz invariant acoustic metric, i.e. $Z^{\mu\nu}=\eta^{\mu\nu}$ and $S_{\mu\nu}=\eta_{\mu\nu}$, in the presence of a gap $C^{2}$ (invariant mass $m^{2}$ in this case). The previous two equations give in this case, the dispersion relation
$\omega^{2}({\bf p})=m^{2}+{\bf p}^{2}\,.$ (44)
Notice that with this, the quantity $\omega^{2}({\bf p})/{\bf p}^{2}$ is divergent for ${\bf p}\to 0$. Instead, the equation
$(N^{0})^{2}=m^{2}+{\bf N}^{2}\,,$ (45)
allows to derive the definition (32) for the ray velocity, which is well-defined at zero momentum. Given the form of the acoustic metric, the ray velocity takes the form:
${\bf v}^{2}=\frac{{\bf N}^{2}}{(N^{0})^{2}}=\frac{{\bf p}^{2}}{m^{2}+{\bf p}^{2}}\leq 1\;,$ (46)
which for this specific case coincides with the group velocity. Equivalently, we find
$N^{0}=\frac{m}{\sqrt{1-{\bf v}^{2}}}\,,$ (47)
that is, ${\bf v}$ plays the role of velocity for a relativistic particle.
4 Bounds from the IR ray velocity
4.1 Conformal superfluid in D=3
We now return our focus to the low-energy ($\omega\ll\mu$) mode $\omega_{-}^{2}({\bf p})$, whose ray velocity will be momentum-dependent on account of the $c_{2,3}$ contributions. By bounding this velocity we can put constraints on these higher dimension coefficients. As discussed in the previous section, this being a gapped mode the proper velocity definition is through the ratio of the components of the ray vector $N^{\mu}$, as in Eq. (32).
From the dispersion relation (22) we can read off the acoustic metric [^6] Actually, we can only read it up to a conformal factor, which for our discussion is irrelevant so we choose $Z^{00}=1$.,
$Z^{\mu\nu}(p)\,p_{\mu}p_{\nu}+m^{2}=0\,,$ (48)
with $Z^{\mu\nu}=diag(-1,\alpha^{2},\alpha^{2})$. An important remark here is that we are treating the contributions from HD operators, responsible for the ${\bf p}$-dependence in $\alpha^{2}({\bf p})$, as perturbative within the regime of validity of the EFT, implicitly giving a condition on $c_{2,3}$ and $p_{\mu}$. This prompts us define a $p_{\mu}$-dependent acoustic metric $Z^{\mu\nu}(p)$, which we then otherwise treat as usual. In particular, its inverse is simply $S_{\mu\nu}=diag(-1,1/\alpha^{2},1/\alpha^{2})$ and trading $p_{\mu}$ for $N^{\mu}=Z^{\mu\nu}(p)p_{\nu}$ can always be unambiguously done perturbatively. Rewriting the EoM in terms of $N^{\mu}$ we get
$-(N^{0})^{2}+\frac{{\bf N}^{2}}{\alpha^{2}}+m^{2}=0\,,$ (49)
which clearly would give the usual ${\bf v}^{2}=\alpha^{2}$ if there was no gap. Instead, in this case we have
$\left(-\frac{1}{{\bf v}^{2}}+\frac{1}{\alpha^{2}}\right){\bf N}^{2}+m^{2}=0\,,$ (50)
where we used Eq. (32). It is useful to define the parameter
$\varepsilon\equiv\frac{{\bf p}\cdot{\bf N}}{m^{2}}=\alpha^{2}\frac{{\bf p}^{2}}{m^{2}}\,,$ (51)
Then we find
$\left(-\frac{1}{{\bf v}^{2}}+\frac{1}{\alpha^{2}}\right)\alpha^{2}\varepsilon+1=0\,,$ (52)
which leads to the following expression for the ray velocity
${\bf v}^{2}=\alpha^{2}\frac{\varepsilon}{1+\varepsilon}\,.$ (53)
As anticipated, the gap induced by the coupling between the scalar and gauge modes slows down the propagation. Notice also that this expression for the ray velocity reduces to the usual phase velocity for momenta well above the gap ($\varepsilon\gg 1$), i.e. ${\bf v}^{2}\simeq\omega_{-}^{2}({\bf p})/{\bf p}^{2}\simeq\alpha^{2}({\bf p})$. In particular, this becomes exact for the front velocity, the limiting velocity at infinite momentum which ultimately determines the support in spacetime of the retarded Green’s function, ${\bf v}^{2}(\infty)=\alpha^{2}(\infty)$.
We now proceed to impose our bound. As discussed, the expected behavior of the ray/phase velocity of a gapless mode in a conformal superfluid in $D=3$ is that it runs from the limiting value $c_{s}^{2}=1/2$ at zero momentum to $1$ at infinite momentum. Here however the gap has the effect of slowing down the mode to a crawl at momenta ${\bf p}^{2}\ll m^{2}$, instead allowing $\alpha^{2}({\bf p})$ in this regime to be practically unbounded. To provide a bound on $\alpha^{2}({\bf p})$, the physically meaningful condition is to demand the ray velocity ${\bf v}^{2}$ to remain below $c_{s}=1/2$ for sufficiently low momenta.
Taking this approach we demand,
${\bf v}^{2}\leq c_{s}^{2}=\frac{1}{2}\,.$ (54)
or, equivalently in terms of $\alpha^{2}({\bf p})$,
$\alpha^{2}({\bf p})\leq\frac{1+\varepsilon}{2\varepsilon}\,.$ (55)
On the other hand, we should not extend this condition too far in momentum, as it is only expected that eventually ${\bf v}^{2}$ will increase towards $1$. So a key question is just how far is far enough. A conservative answer is: far enough to bring the corresponding bound on $\alpha^{2}({\bf p})$ down to $1$, which is in turn the limit that we should impose at very high momenta, when ray velocity is indeed just phase velocity ${\bf v}^{2}\simeq\alpha^{2}({\bf p})$ and should approach $1$. The momentum ${\bf p}_{*}$ at which our bound above (54) becomes $\alpha^{2}({\bf p}_{*})\leq 1$ corresponds to $\varepsilon_{*}=1$, that is $\alpha^{2}({\bf p}_{*}){\bf p}_{*}^{2}=m^{2}$. Assuming moreover that the momentum dependence of $\alpha^{2}({\bf p})$ is always increasing, then we also have $\alpha^{2}({\bf p})\geq 1/2$ for all ${\bf p}$. Therefore, the pivot scale ${\bf p}_{*}$ lies in somewhere in the range
$m^{2}\leq{\bf p}_{*}^{2}\leq 2m^{2}\,.$ (56)
After this point, we instead just consider $\alpha^{2}({\bf p})\leq 1$ for momenta ${\bf p}^{2}\geq{\bf p}_{*}^{2}$.
To summarize our bound, we recall the definition for $\alpha^{2}$ from Eq. (22) and using Eq. (55), this entails
$\tilde{c}_{2}(\varepsilon-1)+\tilde{c}_{3}(\varepsilon+1)\leq\frac{1}{2\varepsilon}\,,$ (57)
with $\varepsilon\in[0,1]$. The optimal choice of $\varepsilon$ actually depends on the values of $\tilde{c}_{2}$ and $\tilde{c}_{3}$. In this form our result can be readily compared Ref. Creminelli:2022onn . For this purpose it is useful to change variables by defining $\frac{(1-\varepsilon)}{(1+\varepsilon)}\equiv 1-\xi^{2}$. For the range $\varepsilon\in[0,1]$ that we consider, the new variable is also $\xi\in[0,1]$, while the bound (57) can be rewritten as
$\tilde{c}_{2}\left(1-\xi^{2}\right)-\tilde{c}_{3}\geq-\frac{\left(1-\frac{\xi^{2}}{2}\right)^{2}}{\xi^{2}}\,.$ (58)
This confirms that our IR approach reproduces correctly the bounds derived in Creminelli:2022onn , see Eq. (58) there.
As a concluding remark, we can consider the limiting case where $b+d\to\infty$, where the vector field decouples. Recalling that $\tilde{c}_{2,3}=c_{2,3}/(b+d)$ and taking this limit in Eq. (57) at fixed $\varepsilon$, gives as a result a trivialization of the bound. In other words, we recover the expected outcome that for the pure scalar theory we are unable to extract a meaningful bound, as discussed in Sec. 2.1. From our current vantage point, this can be simply understood as a consequence of the vanishing of the gap in this limit which closes the window where the bound can effectively be applied.
4.2 Connection with $\langle JJ\rangle$ correlators
In Ref. Creminelli:2022onn analyticity bounds were derived for the conformal superfluid model studied above, by studying a Green’s function associated with the correlator of conserved currents $J^{\mu}$, computed in the case where the gauge field remains non-dynamical. The bounds obtained in this case coincide with our result in Eq. (54), which instead we obtained by studying the low-energy theory as is, allowing the gauge field to be dynamical and therefore kinetically mix with the scalar.
To see the connection between these two approaches, consider the path integral
$Z=\int\mathcal{D}\pi\,e^{i\int d^{3}\!x\,\mathcal{L}(\pi,A_{\mu})}\,.$ (59)
Focusing on the quadratic Lagrangian of Eq. (1), which we now parameterize as
$\mathcal{L}_{2}(\pi,A_{\mu})=\frac{1}{2}\left[\int d^{3}\!x^{\prime}\,\pi(x)G^{-1}_{\pi\pi}(x,x^{\prime})\pi(x^{\prime})+2A_{\mu}(x)J^{\mu}(x)+\int d^{3}\!x^{\prime}A_{\mu}(x)[G^{-1}]^{\mu\nu}(x,x^{\prime})A_{\nu}(x^{\prime})\right]\,,$
we can compute correlators by taking functional derivatives. In particular, taking two functional derivatives with respect to $A_{\mu}$ we get
$\frac{1}{Z}\frac{\delta^{2}Z}{\delta A_{\mu}(x)\delta A_{\nu}(x^{\prime})}\Bigg{|}_{A=0}$ $=$ $\frac{1}{Z}\int\mathcal{D}\pi\,\left[-J^{\mu}(x)J^{\nu}(x^{\prime})+i[G^{-1}]^{\mu\nu}(x,x^{\prime})\right]\,e^{i\int d^{3}\!x\,\mathcal{L}_{2}(\pi,A_{\mu})}$ (61)
$=$ $-\langle 0|T\left\{J^{\mu}(x)J^{\nu}(x^{\prime})\right\}|0\rangle+i[G^{-1}]^{\mu\nu}(x,x^{\prime})\,,$
while, if we go ahead and integrate out the scalar $\pi$, we can compute the Gaussian functional integral explicitly and obtain
$Z=N\,\exp\left({\frac{i}{2}\int_{x,x^{\prime}}\!\!\!\!\!\!\!\!A_{\mu}(x)\left\{[G^{-1}]^{\mu\nu}(x,x^{\prime})-\int_{x_{1},x_{2}}\!\!\!\!\!\!\!\!\!\!\,\tilde{J}^{\mu}(x,x_{1})G_{\pi\pi}(x_{1},x_{2})\tilde{J}^{\nu}(x_{2},x^{\prime})\right\}A_{\nu}(x^{\prime})\,}\right)\,,$
where we rewrote the linear-in-$A_{\mu}$ term in the quadratic Lagrangian as
$A_{\mu}(x)J^{\mu}(x)=\int d^{3}\!x^{\prime}\,A_{\mu}(x)\tilde{J}^{\mu}(x,x^{\prime})\pi(x^{\prime})\,.$ (63)
With this result we can immediately find that the correlator of conserved currents is
$\langle 0|T\left\{J^{\mu}(x)J^{\nu}(x^{\prime})\right\}|0\rangle=i\int d^{3}\!x_{1}d^{3}\!x_{2}\,\tilde{J}^{\mu}(x,x_{1})G_{\pi\pi}(x_{1},x_{2})\tilde{J}^{\nu}(x_{2},x^{\prime})\,.$ (64)
Notice that while this correlator is only conserved up to a contact term $p_{\mu}\langle 0|T\left\{J^{\mu}(x)J^{\nu}(x^{\prime})\right\}|0\rangle\sim D^{\nu}(x)\delta(x-x^{\prime})$, where $D^{\nu}(x)$ is local a differential operator, its retarded version is conserved exactly.
At this point it is also clear that, if the gauge field $A_{\mu}$ where to be promoted to a dynamical field, its effective kinetic term after integrating out the scalar would be the quantity in curly brackets in the right-hand-side of Eq. (4.2), which is also the last expression in Eq. (61). This shows that the $\langle JJ\rangle$ correlator contains the information of the kinetic mixing between the scalar and the gauge field. Indeed, this correlator can potentially shift the pole of the $A_{\mu}$ propagator, which determines the kinematics of the gauge field excitations. In Fourier space
$\det\left([G^{-1}]^{\mu\nu}(\omega,{\bf p})+\hat{J}^{\mu}(-\omega,-{\bf p})G_{\pi\pi}(\omega,{\bf p})\hat{J}^{\nu}(\omega,{\bf p})\right)=0\,,$ (65)
where $\hat{J}^{\mu}(\omega,{\bf p})$ is $-i$ times the Fourier transform of $\tilde{J}^{\nu}(x,x^{\prime})$, with components given explicitly in Eqs. (11) and (12). Upon explicit computation, using a gauge fixing, the above can be seen to give the same condition as Eq. (18). This happens because the eigenmode of the kinetic matrix (10) mixes with $A_{\mu}$ and changes its propagation. Therefore the solution is precisely the mode $\omega_{-}^{2}({\bf p})$ given in Eq. (22), whose ray velocity we studied and where our bounds came from.
4.3 Example of UV completion
In this Section we outline a specific example of a UV completion of the conformal superfluid in $D=3$ in the absence of the gauge field, consisting of a conformal complex scalar field, as also exhibited in Ref. Creminelli:2022onn . This provides an explicit realization of the momentum dependence of the ray/phase velocity of the gapless mode, which transitions from $1/2$ at zero momentum to $1$ at momenta much larger than the scale of symmetry breaking $\mu$.
The Lagrangian for this UV-complete model is given by Badel:2019khk ,
$S_{\mathrm{UV}}$ $=$ $\sqrt{-g}\left(-|\partial\phi|^{2}-\lambda|\phi|^{6}-\frac{1}{8}R|\phi|^{2}\right)$ (66)
$=$ $\sqrt{-g}\left(-(\partial\rho)^{2}-\rho^{2}(\partial\theta)^{2}-\lambda\rho^{6}-\frac{1}{8}R\rho^{2}\right)\,,$
where $\phi=\rho\,e^{i\theta}$ and $\lambda>0$. Upon integrating out the modulus field $\rho$ at tree level, the EFT action from Eq. (1) is obtained with coefficients Creminelli:2022onn
$c_{1}=\frac{4}{\sqrt{3\lambda}}\,,\qquad c_{2}=\frac{1}{8\sqrt{3\lambda}}\,,\quad c_{3}=b=d=0\,,$ (67)
plus higher-dimensional operators. Coupling to an external gauge field $A_{\mu}$ can be done by introducing a gauge-covariant derivative $\partial_{\mu}\to\nabla_{\mu}$. Expanding around a state with chemical potential $\mu$, $\theta=\mu t+\frac{\pi_{c}}{\sqrt{2}\bar{\rho}}$ and $\rho=\bar{\rho}+\frac{\delta\rho_{c}}{\sqrt{2}}$, where $\bar{\rho}\simeq\sqrt{\mu}/(3\lambda)^{1/4}$, the quadratic lagrangian for the fluctuations $\pi_{c}$ and $\delta\rho_{c}$ has the following kinetic matrix in Fourier space
$\mathcal{K}_{IJ}(\omega,{\bf p})=\begin{pmatrix}\omega^{2}-{\bf p}^{2}&2i\mu\omega\\ -2i\mu\omega&\quad\omega^{2}-{\bf p}^{2}-4\mu^{2}\\ \end{pmatrix}\,.$ (68)
The corresponding eigenmodes $\omega({\bf p})$ are obtained from $\det(\mathcal{K})=0$,
$\omega_{\phi\pm}^{2}({\bf p})={\bf p}^{2}+4\mu^{2}\left(1\pm\sqrt{1+\frac{{\bf p}^{2}}{4\mu^{2}}}\right)\,.$ (69)
The minus branch gives rise to a gapless mode, the superfluid phonon, while the plus branch describes a masive mode with $M^{2}=8\mu^{2}$. Within the regime of validity of the EFT, only the former can be excited. It is instructive to examine its dispersion relation in both the regimes of low and high momentum $|{\bf p}|$ with respect to $\mu$,
$\omega_{\phi-}^{2}({\bf p})\simeq\begin{cases}\frac{{\bf p}^{2}}{2}&\qquad|{\bf p}|\ll\mu\,,\\ {\bf p}^{2}&\qquad|{\bf p}|\gg\mu\,.\end{cases}$ (70)
The low-energy side of this should be compared to Eq. (8), i.e. the gapless mode of the EFT without any dynamical gauge field.
Since this is a gapless mode, the ray and phase velocities coincide and can be read directly from ${\bf v}^{2}=\omega_{\phi-}^{2}({\bf p})/{\bf p}^{2}$. The above exhibits how the ray/phase velocity ${\bf v}^{2}$ of this mode interpolates between the soundspeed of a conformal superfluid in $D=3$, $c_{s}^{2}=1/2$, at low energies when the Lorentz invariance and the conformal invariance are spontaneously broken by the finite chemical potential $\mu$, and the speed of light $c^{2}=1$ in the unbroken phase at high energies.
5 Discussion
In this work we have addressed the question of what is the IR characterization of analyticity/positivity bounds in theories with spontaneously broken Lorentz symmetry. We have derived bounds in terms of the ray velocity of low energy modes in the EFT, capturing exactly the bounds derived from analyticity. We studied the context of an EFT for a conformal superfluid in $D=3$, presented in Sec. 2, studying an IR model in which the $U(1)$ symmetry group is gauged and the superfluid is coupled to a gauge vector field. In this model the EFT has a mass gap, meaning that the ray velocity goes to zero for vanishing momenta. In Sec. 4 we posed IR bounds by requiring that the ray velocity doesn’t grow too fast for low momenta. In the specific, we have imposed that the ray velocity does not grow beyond the speed of gapless excitations, $c_{s}^{2}=1/2$, for momenta below the mass gap, see Eq. (54).
Technically, the similarity between our bound and that of Creminelli:2022onn can be traced back to the close connection between correlators of conserved currents and propagators of the low energy modes, which we discussed in Sec. 4.2.
In our approach, we consider the vector field in the IR theory to be dynamical. Conservatively, this EFT can be thought of as an auxiliary IR system from which to extract the information contained in analyticity bounds. It would be interesting to investigate whether this system can also be considered physical, as a stable low energy phase of a conformally invariant UV theory. (See e.g. Kaplan:2023fbl ; Seiberg:2024wgj ; Berezhiani:2024boz for recent progress on gauge theories at finite charge density.) In this regard, it would be non-trivial to show that one can have a phase with vanishing electric field, even using a background charge density. Another interesting point is that the EFT we study only has gapped modes. It will be useful to explain this either in terms of analogy with plasma physics, or in terms of the algebra of broken generators, see e.g. Nicolis:2012vf .
In conclusion, we have learned that the requirement that modes below the mass gap treshold should be slower than the gapless modes leads to meaningful bounds on the IR coefficients. This reproduces the known analyticity bounds. Our work moves in the direction of developing a simple way to infer the UV properties of nature from IR quantities, even when the systems observed cannot be described within the most conservative assumptions.
As a direct extension of our work one might consider bounds derived for conformal superfluids in Ref. Creminelli:2022onn , for other higher-order operators that we have not studied here. The analyticity bounds for these operators were derived studying two-point correlators of the stress-energy tensor. On the IR side, learning from our results one may expect that the coupling to gravity will be important in this case. Considering the propagation of modes which mix the conformal scalar and gravitational excitations may provide the necessary handle to bound the coefficients of these irrelevant operators from the IR.
In the future it would also be interesting to turn our attention towards other domains where these ideas might be applicable. Indeed, while the conformal superfluid EFT that we studied here is interesting in its own right, it can also be considered as a proxy for more complicated theories in which Lorentz invariance is spontaneously broken. Such is the case of cosmological theories of modified gravity, our original motivation.
Acknowledgements.
The authors would like to thank Gabriel Cuomo, Ignacy Sawicki, Georg Trenkler and Enrico Trincherini for useful discussions, and Alessandro Podo for discussions and comments on the earlier version of this work. The work of F.S. was supported by the Simons Investigator Award No. 827042 (P.I.: Surjeet Rajendran) and the Simons Investigator Award No. 144924 (P.I.: David E. Kaplan). The work of L.G.T. was supported by European Union (Grant No. 101063210).